Before prorating the manufacturing overhead costs at the end of 2016, the Cost of Goods Sold and Finished Goods Inventory had applied overhead costs of $58,700 and $32,000 in them, respectively. There was no Work-in-Process at the beginning or end of 2016. During the year, manufacturing overhead costs of $86,000 were actually incurred. The balance in the Applied Manufacturing Overhead was $90,700 at the end of 2016. If the under- or overapplied overhead is prorated between Cost of Goods Sold and the inventory accounts, how much will be the Cost of Goods Sold after the proration?

Multiple Choice

  • $57,042.

  • $59,617.

  • $60,358.

  • $55,658.
Answer #1
Overhead applied Proration of overhead applied Over applied overhead to be deducted Overhead after proration
Cost of goods old 58,700 58,700/90,700 = 64.71885336% 4,700 x 64.71885336% = $3,042 58,700 - 3,042 = $55,658
Finished goods 32,000 32,000/90,700 = 35.28114664% 4,700 x 35.28114664% = $1,658 32,000 - 1,658 = $30,342
Total 90,700 100% $86,000

Actual overhead = $86,000

Applied overhead = $90,700

Over applied overhead = Applied overhead - Actual overhead

= 90,700 - 86,000

= $4,700

Cost of goods sold after proration = 55,658

Fourth option is the correct option

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.